Last month, you lent a work colleague $5000 to cover some overdue bills. He agreed to pay you in 1 month with interest at 2% for the month, thus owing you $5100. Today, when the repayment is due, he asked you to extend the loan for another month and he would pay you the $5100 next month. In the meantime, you have had the offer to invest as much as you wish in an oil-well venture that is expected to pay 20% per year and a hot new IT stock that is estimated to return 48% the first year. If you let your colleague have another month, what is the opportunity cost of your decision? (Note: Express your answer in dollar and percentage amounts.)

The opportunity cost is $  .

The opportunity cost in percentage is  %
Answer #1

The opportunity cost is $100

Explanation: The opportunity cost is the value of the next-highest-valued alternative use of that resource. Thus opportunity cost will be new IT stock that is estimated to return 48% (=5000 * 48% * 1/12 = 200) will be the opportunity cost instead of an investment in an oil-well venture which will yield lesser return per year. Thus opportunity cost equals 200 - 100 = 100

Opportunity percentage : $100 / 100 * 100 = 100%